A 2µF capacitor is charged to 10 V, and then discharged through a 1.0 MΩ resistor.


a) What is the time constant of the discharge circuit? (Answer: 2 s)


b) What is the time taken for the potential difference across the capacitor to drop to 5 V? (Answer: 1.39


s)


1
Expert's answer
2022-01-25T07:09:04-0500

"\tau=RC=1×10^6×2×10^{-6}"


"=2seconds"


"V=E(e^{\frac{-t}{RC}})"


"5=10(e^\frac{-t}{2})"


Solving for t,


t=1.3863 seconds